# Stealthy Yet Effective: Distribution-Preserving Backdoor Attacks on Graph Classification

### v1

@@ Abstract
Graph Neural Networks (GNNs) have demonstrated strong performance across tasks such as node classification, link prediction, and graph classification, but remain vulnerable to backdoor attacks that implant imperceptible triggers during training to control predictions. While node-level attacks exploit local message passing, graph-level attacks face the harder challenge of manipulating global representations while maintaining stealth. We identify two main sources of anomaly in existing graph classification backdoor methods: structural deviation from rare subgraph triggers and semantic deviation caused by label flipping, both of which make poisoned graphs easily detectable by anomaly detection models. To address this, we propose DPSBA, a clean-label backdoor framework that learns in-distribution triggers via adversarial training guided by anomaly-aware discriminators. DPSBA effectively suppresses both structural and semantic anomalies, achieving high attack success while significantly improving stealth. Extensive experiments on real-world datasets validate that DPSBA achieves a superior balance between effectiveness and detectability compared to state-of-the-art baselines.
@@ 1 Introduction
Graph Neural Networks (GNNs) are foundational models for learning on graph-structured data, achieving strong performance in tasks like node classification, link prediction, and graph classification [1]. Through message passing, GNNs capture structural and feature dependencies, enabling applications in social networks [2], recommender systems [3], and molecular analysis [4]. As GNNs are increasingly adopted in real-world systems, their security has drawn growing attention. Among various threats, backdoor attacks, where imperceptible triggers are embedded in training data to manipulate predictions, pose a serious risk due to their stealth [5, 6, 7]. While extensively studied in vision and NLP, backdoor vulnerabilities in graph learning remain under exploration.
Most existing graph backdoor studies focus on node classification, where local triggers are injected to misclassify specific nodes [8, 9]. These attacks leverage GNN locality to propagate triggers through neighborhoods while remaining hard to detect. Link prediction attacks similarly operate in a local context, targeting the presence or absence of edges between node pairs [10]. In contrast, graph classification poses a fundamentally different and more complex challenge. Here, the attacker must influence the global semantics of an entire graph rather than a single node or edge. This requires triggers to manipulate the full-graph embedding, often through larger or structurally rare subgraphs.
@@
Figure 1: Anomaly score distributions of clean vs. backdoor graphs on the AIDS dataset. Anomaly scores are computed via SIGNET [11] for backdoor samples generated by ER-B [12], GTA [13], and Motif [14]. In all cases, backdoor graphs (yellow) shift significantly rightward compared to clean graphs (blue), revealing strong distributional deviation. This underscores the detectability of prior triggers and the need for distribution-preserving attacks.
Recent backdoor attacks on graph classification attempt to address global manipulation challenges by injecting structurally distinct subgraphs, such as rare motifs [14]. However, these approaches often introduce obvious out-of-distribution (OOD) artifacts, which significantly compromise stealth and limit their practicality in real-world settings. As shown in Figure 1, there exists a clear distribution gap between clean and backdoor graphs, making current methods highly detectable by outlier detection models. Although Zhang et al. [15] similarly observe OOD issues in node classification and propose a distribution-preserving strategy, their approach relies on local perturbations that affect only a single node’s embedding. In contrast, graph classification requires modifying the global message-passing process to influence the entire graph representation, making distribution preservation significantly more challenging (see Appendix C.1 for theoretical justification).
@@ 3.1.1 Hard Sample Selection
In the clean-label setting, directly selecting target-class samples avoids semantic inconsistency but may reduce the effectiveness of the attack. To counteract this, DPSBA strategically mines hard samples, i.e., samples from the target class that the model finds uncertain. This design is inspired by Gao et al. [16], who show that hard samples can amplify the effectiveness of backdoor attacks by being more susceptible to manipulation. In particular, low-confidence samples, those near the model’s decision boundary, require smaller perturbations to shift predictions, and are thus better candidates for subtle backdoor injection. We use a clean surrogate model $f_{\theta}$ to measure the confidence score of a graph $G$ with respect to the target label $y_{t}$ as:
@@ 3.1.2 Trigger Location Selection
After selecting the hard target-class samples for poisoning, DPSBA determines where in each graph to inject the trigger. Existing studies adopt various strategies to identify influential nodes, such as using centrality metrics [14] or extracting interpretable subgraphs [17, 18]. While interpretation-based methods offer precision, they often incur prohibitive computational costs. To strike a balance between scalability and informativeness, DPSBA adopts a two-stage selection process. First, nodes with high degree centrality are pre-selected as candidates. Specifically, for a graph $G$ with $N$ nodes, we compute the normalized degree centrality of each node $v$ as $\text{deg}(v)/(N-1)$ and retain the top-$k$ nodes (with $k$ typically set to $2M$, where $M$ is the trigger size). Next, among these candidates, we identify the $M$ most influential nodes using a surrogate-based ablation approach. For each node $v$ in the candidate set $\mathcal{N}_{can}$, we evaluate its contribution to model prediction by measuring the change in output after removing it:
@@ 3.1.3 Trigger Generation and Injection
To enable the insertion of stealthy and effective triggers, generative approaches offer superior flexibility and control over random, interpretability-based, or search-based strategies. Inspired by GTA [13], DPSBA introduces two lightweight yet expressive modules, i.e., a topology generator and a feature generator, that cooperatively define the structure and node attributes of the injected subgraph. These generators are trained in a separate phase to adaptively fit the data distribution.
@@
The topology generator is implemented as a multilayer perceptron (MLP) that transforms the adjacency matrix $\mathbf{H}$ of the target injection region into a learnable soft structure: $\mathbf{H^{\prime}}=\sigma(W_{1}\mathbf{H}+b_{1}),$ where $W_{1}$ and $b_{1}$ are learnable parameters (for the convenience of subsequent discussion, they will be simply referred to as $\omega_{t}$}) and $\sigma(\cdot)$ is a non-linear activation function. For undirected graphs, the adjacency matrix is symmetrized as $\mathbf{A}=\frac{1}{2}\left(\mathbf{H^{\prime}}+\mathbf{H^{\prime}}^{\top}\right)$. To accommodate the discrete nature of graph topology, we adopt a binarization mechanism inspired by binarized neural networks [19], where the final adjacency matrix is given by: $\mathbf{A}_{binary}=\mathbb{I}(\mathbf{A}>0.5).$ Binarization is applied only during forward propagation to preserve gradient flow during backpropagation. The resulting binary matrix defines the trigger’s structural connectivity.
@@ 4.1.1 Datasets and Evaluation Metrics
We evaluate DPSBA on four real-world graph classification datasets from the TUDataset benchmark [20]: PROTEINS_full [21] ( protein graphs for function prediction), AIDS [22] (molecular graphs related to AIDS research), FRANKENSTEIN [23] (a compound property dataset combining BURS and MNIST features), and ENZYMES [24, 25] (a 6-class biomolecular classification task). For each dataset, we designate the minority class as the attack target to simulate a realistic low-frequency scenario. The statistics for the datasets are summarized in Table E7. We adopt three metrics: ASR (Attack Success Rate) for attack effectiveness, CAD (Clean Accuracy Drop) to assess model utility degradation, and AUC (Area Under the ROC Curve) to quantify anomaly detectability via outlier models. Detailed metric definitions are deferred to Appendix E.2.
@@ 4.1.2 Baselines
We compare DPSBA with five representative graph-level backdoor attack methods. ER-B [12] generates universal triggers using the Erdős–Rényi random graph model. LIA [17] modifies the connections of low-importance nodes, as identified by GNN explanation techniques, with fixed trigger structures. GTA [13] employs a bi-level optimization strategy to train a subgraph generator that adapts triggers per graph. Motif [14] selects low-frequency motifs from the dataset as effective but easily detectable triggers. Motif-S is a stealthier variant using the high-frequency M41 motif, reducing anomaly scores while slightly sacrificing attack strength. Following GTA [13], we split each dataset into 50% training and 50% test sets, with 5% of the training data poisoned. To ensure fair comparison, all methods adopt a fixed trigger size of 4. Both the topology and feature generators are trained for 20 epochs per stage over 3 iterations with a learning rate of 0.001, using early stopping [26]. For baselines, we use the best hyperparameters reported in their original papers.
@@ 4.1.3 Graph Classification Models and Anomaly Detection Algorithms
Following prior work on graph backdoor attacks [14], we evaluate DPSBA on three widely used graph classifiers: GCN [27], GIN [28], and SAGPool [29]. These models are used both as attack targets and surrogates to test generalizability across architectures. Clean accuracy results are reported in Table E8. For anomaly detection, we adopt SIGNET [11], identified as the most effective method in a recent benchmark study [30], to assess the stealthiness of injected backdoor samples.
@@ 4.2.1 Effectiveness and Stealthiness
We evaluate DPSBA on three datasets against six baselines in terms of attack effectiveness (ASR), model performance drop (CAD), and anomaly detectability (AUC), as shown in Table 1. DPSBA consistently achieves optimal or near-optimal ASR across datasets and models, verifying its strong attack capability. This is credited to its adaptive trigger generation guided by confident hard samples and structurally important nodes. In terms of CAD, DPSBA maintains minimal accuracy degradation (<5%), indicating high stealth at the model level. For distribution-level stealth, DPSBA significantly reduces AUC, i.e., maintaining values around 70%, demonstrating strong resistance to detection by statistical outlier models. We further observe that ASR and AUC fluctuate significantly across datasets due to inherent differences in graph structure and anomaly sensitivity. For datasets particularly sensitive to attribute or structural perturbations (e.g., FRANKENSTEIN), even minor modifications can induce high ASR yet large anomaly scores. In such cases, we recommend an early stopping strategy, namely terminating generator updates once ASR saturates or marginal gains diminish, to enhance stealth. For instance, on FRANKENSTEIN, where attribute anomalies are highly detectable (GTA achieves >95% ASR but with high AUC), DPSBA disables the feature generator early, achieving near-perfect ASR solely via structural triggers. Across all benchmarks, DPSBA achieves the high ASR of Motif while matching the stealth of Motif-S, validating its ability to balance effectiveness and detectability. For multi-class graph classification tasks, DPSBA consistently outperforms all baselines in terms of both attack success rate (ASR) and stealth (CAD, AUC) across three surrogate models. Notably, most baselines (except GTA) fail on this dataset due to high attribute and strucutre variability, underscoring the advantage of our distribution-preserving and adaptive design.
@@ 4.2.3 Anomaly Stealth Analysis
To show the anomaly stealthiness of DPSBA, we first visualize the anomaly score distributions of clean and backdoor graphs on all datasets in Figure 3. The distributions show strong overlap, indicating that DPSBA-generated triggers exhibit minimal anomaly and are hard to distinguish from clean samples. This is largely attributed to the use of clean-label constraints and adaptive optimization. To further assess robustness across detectors, we evaluate DPSBA under multiple anomaly detection algorithms, including OCGIN [31] and GLocalKD [32], with results summarized in Table 3. DPSBA consistently achieves low AUCs and, notably, GLocalKD fails to detect most triggers, confirming the strong stealth of DPSBA across detection methods. Beyond anomaly detection, we examine randomized subsampling [12, 13], a standard defense strategy that randomly removes subgraphs during training. As shown in Figure 4(a), DPSBA maintains high ASR despite this defense, thanks to its compact trigger size and dual embedding in structure and features, making the trigger difficult to eliminate via sampling.
@@
4.3 Impact of the Loss Weights $\mathbf{\alpha}$} and $\mathbf{\beta}$}
Figure 5: Parametric analysis of $\alpha$ and $\beta$
@@ 6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62302333, No. 92370111, No. 62272340, No. 62422210 and No. 62276187) and the Open Research Fund from Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) (No.GML-KF-24-16).
@@
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.

### v2

@@ Abstract
Graph Neural Networks (GNNs) have demonstrated strong performance across tasks such as node classification, link prediction, and graph classification, but remain vulnerable to backdoor attacks that implant imperceptible triggers during training to control predictions. While node-level attacks exploit local message passing, graph-level attacks face the harder challenge of manipulating global representations while maintaining stealth. We identify two main sources of anomaly in existing graph classification backdoor methods: structural deviation from rare subgraph triggers and semantic deviation caused by label flipping, both of which make poisoned graphs easily detectable by anomaly detection models. To address this, we propose DPSBA, a clean-label backdoor framework that learns in-distribution triggers via adversarial training guided by anomaly-aware discriminators. DPSBA effectively suppresses both structural and semantic anomalies, achieving high attack success while significantly improving stealth. Extensive experiments on real-world datasets validate that DPSBA achieves a superior balance between effectiveness and detectability compared to state-of-the-art baselines. The code is available at https://github.com/TheCoderOfs/DPSBA.
@@ 1 Introduction
Graph Neural Networks (GNNs) are foundational models for learning on graph-structured data, achieving strong performance in tasks like node classification, link prediction, and graph classification [wu2020comprehensive]. Through message passing, GNNs capture structural and feature dependencies, enabling applications in social networks [fan2019graph], recommender systems [gao2022graph], and molecular analysis [wieder2020compact]. As GNNs are increasingly adopted in real-world systems, their security has drawn growing attention. Among various threats, backdoor attacks, where imperceptible triggers are embedded [Liu2024EfficientUG] in training data to manipulate predictions, pose a serious risk due to their stealth [zhang2023graph, dai2023unnoticeable, FENG2024106668]. While extensively studied in vision and NLP, backdoor vulnerabilities in graph learning remain under exploration.
Most existing graph backdoor studies focus on node classification, where local triggers are injected to misclassify specific nodes [chen2023feature, yang2023percba]. These attacks leverage GNN locality to propagate triggers through neighborhoods while remaining hard to detect. Link prediction attacks similarly operate in a local context, targeting the presence or absence of edges between node pairs [dai2024backdoor]. In contrast, graph classification poses a fundamentally different and more complex challenge. Here, the attacker must influence the global semantics of an entire graph rather than a single node or edge. This requires triggers to manipulate the full-graph embedding, often through larger or structurally rare subgraphs.
@@
Figure 1: Anomaly score distributions of clean vs. backdoor graphs on the AIDS dataset. Anomaly scores are computed via SIGNET [liu2023towards] for backdoor samples generated by ER-B [zhang2021backdoor], GTA [xi2021graph], and Motif [zheng2023motif]. In all cases, backdoor graphs (yellow) shift significantly rightward compared to clean graphs (blue), revealing strong distributional deviation. This underscores the detectability of prior triggers and the need for distribution-preserving attacks.
Recent backdoor attacks on graph classification attempt to address global manipulation challenges by injecting structurally distinct subgraphs, such as rare motifs [zheng2023motif]. However, these approaches often introduce obvious out-of-distribution (OOD) artifacts, which significantly compromise stealth and limit their practicality in real-world settings. As shown in Figure 1, there exists a clear distribution gap between clean and backdoor graphs, making current methods highly detectable by outlier detection models. Although Zhang et al. [10.1145/3637528.3671910] similarly observe OOD issues in node classification and propose a distribution-preserving strategy, their approach relies on local perturbations that affect only a single node’s embedding. In contrast, graph classification requires modifying the global message-passing process to influence the entire graph representation, making distribution preservation significantly more challenging (see Appendix C.1 for theoretical justification).
@@ 3.1.1 Hard Sample Selection
In the clean-label setting, directly selecting target-class samples avoids semantic inconsistency but may reduce the effectiveness of the attack. To counteract this, DPSBA strategically mines hard samples, i.e., samples from the target class that the model finds uncertain. This design is inspired by Gao et al. [2023Not], who show that hard samples can amplify the effectiveness of backdoor attacks by being more susceptible to manipulation. In particular, low-confidence samples, those near the model’s decision boundary, require smaller perturbations to shift predictions, and are thus better candidates for subtle backdoor injection. We use a clean surrogate model $f_{\theta}$ to measure the confidence score of a graph $G$ with respect to the target label $y_{t}$ as:
@@ 3.1.2 Trigger Location Selection
After selecting the hard target-class samples for poisoning, DPSBA determines where in each graph to inject the trigger. Existing studies adopt various strategies to identify influential nodes, such as using centrality metrics [zheng2023motif] or extracting interpretable subgraphs [xu2021explainability, wang2024explanatory]. While interpretation-based methods offer precision, they often incur prohibitive computational costs. To strike a balance between scalability and informativeness, DPSBA adopts a two-stage selection process. First, nodes with high degree centrality are pre-selected as candidates. Specifically, for a graph $G$ with $N$ nodes, we compute the normalized degree centrality of each node $v$ as $\text{deg}(v)/(N-1)$ and retain the top-$k$ nodes (with $k$ typically set to $2M$, where $M$ is the trigger size). Next, among these candidates, we identify the $M$ most influential nodes using a surrogate-based ablation approach. For each node $v$ in the candidate set $\mathcal{N}_{can}$, we evaluate its contribution to model prediction by measuring the change in output after removing it:
@@ 3.1.3 Trigger Generation and Injection
To enable the insertion of stealthy and effective triggers, generative approaches offer superior flexibility and control over random, interpretability-based, or search-based strategies. Inspired by GTA [xi2021graph], DPSBA introduces two lightweight yet expressive modules, i.e., a topology generator and a feature generator, that cooperatively define the structure and node attributes of the injected subgraph. These generators are trained in a separate phase to adaptively fit the data distribution.
@@
The topology generator is implemented as a multilayer perceptron (MLP) that transforms the adjacency matrix $\mathbf{H}$ of the target injection region into a learnable soft structure: $\mathbf{H^{\prime}}=\sigma(W_{1}\mathbf{H}+b_{1}),$ where $W_{1}$ and $b_{1}$ are learnable parameters (for the convenience of subsequent discussion, they will be simply referred to as $\omega_{t}$}) and $\sigma(\cdot)$ is a non-linear activation function. For undirected graphs, the adjacency matrix is symmetrized as $\mathbf{A}=\frac{1}{2}\left(\mathbf{H^{\prime}}+\mathbf{H^{\prime}}^{\top}\right)$. To accommodate the discrete nature of graph topology, we adopt a binarization mechanism inspired by binarized neural networks [hubara2016binarized], where the final adjacency matrix is given by: $\mathbf{A}_{binary}=\mathbb{I}(\mathbf{A}>0.5).$ Binarization is applied only during forward propagation to preserve gradient flow during backpropagation. The resulting binary matrix defines the trigger’s structural connectivity.
@@ 4.1.1 Datasets and Evaluation Metrics
We evaluate DPSBA on four real-world graph classification datasets from the TUDataset benchmark [Morris+2020]: PROTEINS_full [borgwardt2005protein] ( protein graphs for function prediction), AIDS [rossi2015network] (molecular graphs related to AIDS research), FRANKENSTEIN [orsini2015graph] (a compound property dataset combining BURS and MNIST features), and ENZYMES [10.1093/bioinformatics/bti1007, schomburg2004brenda] (a 6-class biomolecular classification task). For each dataset, we designate the minority class as the attack target to simulate a realistic low-frequency scenario. The statistics for the datasets are summarized in Table E7. We adopt three metrics: ASR (Attack Success Rate) for attack effectiveness, CAD (Clean Accuracy Drop) to assess model utility degradation, and AUC (Area Under the ROC Curve) to quantify anomaly detectability via outlier models. Detailed metric definitions are deferred to Appendix E.2.
@@ 4.1.2 Baselines
We compare DPSBA with five representative graph-level backdoor attack methods. ER-B [zhang2021backdoor] generates universal triggers using the Erdős–Rényi random graph model. LIA [xu2021explainability] modifies the connections of low-importance nodes, as identified by GNN explanation techniques, with fixed trigger structures. GTA [xi2021graph] employs a bi-level optimization strategy to train a subgraph generator that adapts triggers per graph. Motif [zheng2023motif] selects low-frequency motifs from the dataset as effective but easily detectable triggers. Motif-S is a stealthier variant using the high-frequency M41 motif, reducing anomaly scores while slightly sacrificing attack strength. Following GTA [xi2021graph], we split each dataset into 50% training and 50% test sets, with 5% of the training data poisoned. To ensure fair comparison, all methods adopt a fixed trigger size of 4. Both the topology and feature generators are trained for 20 epochs per stage over 3 iterations with a learning rate of 0.001, using early stopping [yao2007early]. For baselines, we use the best hyperparameters reported in their original papers.
@@ 4.1.3 Graph Classification Models and Anomaly Detection Algorithms
Following prior work on graph backdoor attacks [zheng2023motif], we evaluate DPSBA on three widely used graph classifiers: GCN [kipf2017semisupervised], GIN [xu2018how], and SAGPool [lee2019self]. These models are used both as attack targets and surrogates to test generalizability across architectures. Clean accuracy results are reported in Table E8. For anomaly detection, we adopt SIGNET [liu2023towards], identified as the most effective method in a recent benchmark study [wang2025unifying], to assess the stealthiness of injected backdoor samples.
@@ 4.2.1 Effectiveness and Stealthiness
We evaluate DPSBA on four datasets against six baselines in terms of attack effectiveness (ASR), model performance drop (CAD), and anomaly detectability (AUC), as shown in Table 1. DPSBA consistently achieves optimal or near-optimal ASR across datasets and models, verifying its strong attack capability. This is credited to its adaptive trigger generation guided by confident hard samples and structurally important nodes. In terms of CAD, DPSBA maintains minimal accuracy degradation (<5%), indicating high stealth at the model level. For distribution-level stealth, DPSBA significantly reduces AUC, i.e., maintaining values around 70%, demonstrating strong resistance to detection by statistical outlier models. We further observe that ASR and AUC fluctuate significantly across datasets due to inherent differences in graph structure and anomaly sensitivity. For datasets particularly sensitive to attribute or structural perturbations (e.g., FRANKENSTEIN), even minor modifications can induce high ASR yet large anomaly scores. In such cases, we recommend an early stopping strategy, namely terminating generator updates once ASR saturates or marginal gains diminish, to enhance stealth. For instance, on FRANKENSTEIN, where attribute anomalies are highly detectable (GTA achieves >95% ASR but with high AUC), DPSBA disables the feature generator early, achieving near-perfect ASR solely via structural triggers. Across all benchmarks, DPSBA achieves the high ASR of Motif while matching the stealth of Motif-S, validating its ability to balance effectiveness and detectability. For multi-class graph classification tasks, DPSBA consistently outperforms all baselines in terms of both attack success rate (ASR) and stealth (CAD, AUC) across three surrogate models. Notably, most baselines (except GTA) fail on this dataset due to high attribute and strucutre variability, underscoring the advantage of our distribution-preserving and adaptive design.
@@ 4.2.3 Anomaly Stealth Analysis
To show the anomaly stealthiness of DPSBA, we first visualize the anomaly score distributions of clean and backdoor graphs on all datasets in Figure 3. The distributions show strong overlap, indicating that DPSBA-generated triggers exhibit minimal anomaly and are hard to distinguish from clean samples. This is largely attributed to the use of clean-label constraints and adaptive optimization. To further assess robustness across detectors, we evaluate DPSBA under multiple anomaly detection algorithms, including OCGIN [zhao2023using] and GLocalKD [ma2022deep], with results summarized in Table 3. DPSBA consistently achieves low AUCs and, notably, GLocalKD fails to detect most triggers, confirming the strong stealth of DPSBA across detection methods. Beyond anomaly detection, we examine randomized subsampling [zhang2021backdoor, xi2021graph], a standard defense strategy that randomly removes subgraphs during training. As shown in Figure 4(a), DPSBA maintains high ASR despite this defense, thanks to its compact trigger size and dual embedding in structure and features, making the trigger difficult to eliminate via sampling.
@@ 4.3 Ablation Study
4.4 Impact of the Loss Weights $\mathbf{\alpha}$} and $\mathbf{\beta}$}
Figure 5: Parametric analysis of $\alpha$ and $\beta$
@@ 6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62302333, No. 92370111, No. 62272340, No. 62422210, No. 62276187 and No. 62502065) , the Open Research Fund from Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) (No. GML-KF-24-16), and Hebei Natural Science Foundation (No. F2024202047).